Sodium bicarbonate reacts with citric acid Alka Seltzer tablets when added to water in the following balanced reaction: 3NaHCO3 + C6H8O7 —> Na3C6H5O7 + 3CO2 + 3H2O a. What mass of carbon dioxide can be made from the reaction of 22.7 grams of NaHCO3 with 17.5 grams of C6H8O7 b. What is the limiting reactant ?


1
Expert's answer
2021-03-15T09:08:49-0400

Moles of NaHCO3 = 22.7/84

= 0.270 mol

Moles of C6H8O7 = 17.5/192

= 0.0911 mol

Limiting reagent = [NaHCO3]

Mass of CO2 = 3×44×22.7/3×84

= [11.9] g